Chapter 7 -  BB exercise  No. 2 - Practice Question no 1

MSU makes blue jean jackets.  The standard costs per jacket are as follows:

                                            Budgeted Qty          Budgeted Price

Direct Material                                   .20                               $ 60.00 per yard

Direct Labor                                        .10                               $ 30.00 per hour

MSU produced and sold 10,000 jackets using 1,000 yards of direct materials at a cost of $ 64.00 per yard and 500 hours of direct labor at a wage of $ 30.50 per hour.

What is the direct manufacturing labor efficiency variance? Also, is the variance favorable or unfavorable?
Answer #1

Labor efficiency variance = (SH-AH) * SR

= (10,000*0.10 - 500) * 30

= (1000-500) * 30

= 1500 Favourable